The composition of a flake of paint found on the c
Answer #1

(a). Paint from clothes -

x x - \bar{x} (x - \bar{x})2
4.0 - 0.3 0.09
4.2 - 0.1 0.01
4.4 0.1 0.01
4.6 0.3 0.09
\Sigma 17.2 0.2

Total number of data points, N = 4

Mean value of %T of paint from clothes:

\bar{x} = \Sigma x / N

= (4.0 + 4.2 + 4.4 + 4.6) / 4

= 4.3

Standard deviation, S.D = \sqrt{\Sigma (x - \bar{x})^2) / N}

= \sqrt{0.2 / 4}

= \sqrt{0.05}

= 0.22

Relative standard deviation = S.D / \bar{x} * 100

= 0.22 / 4.3 * 100

= 5.1 %

Paint from Car -

x x - \bar{x} (x - \bar{x})2
4.5 - 0.54 0.2916
5.3 0.26 0.0676
5.5 0.46 0.2116
5.0 - 0.04 0.0016
4.9 - 0.14 0.0196
\Sigma 25.2 0.592

Mean value of %T of paint from car:

\bar{x} = \Sigma x / N

= 25.2 / 5

= 5.0

Standard deviation, S.D = \sqrt{\Sigma (x - \bar{x})^2) / N}

= \sqrt{0.592 / 5}

= 0.34

Relative standard deviation = S.D / \bar{x} * 100

= 0.34 / 5.0 * 100

= 6.8%

(b). S1 = 0.22 , N1 = 4

S2 = 0.34,   N2 = 5

Pooled standard deviation,

(ni + n2-2)

= \sqrt{}[(4-1)*0.222 + (5-1)*0.342] / (4+5-2)

= \sqrt{}[( 3*0.0484 + 4*0.1156 ) / 7]

= \sqrt{}(0.1452 + 0.4624) / 7

= \sqrt{}0.0868

= 0.29